The sum of jack and Susan’s age is 35. The difference of their age is 27 how old is each person?
Answer by ankor@dixie-net.com(22740)

The sum of jack and Susan’s age is 35. The difference of their age is 27 how old is each person?
:
j + s = 35
j - s = 27
--------------addition eliminates s, find j
2j + 0 = 62
j = 62/2
j = 31 is Jack age
then obviously, 4 is Susan's age